Operations Report – May 4th

Crew 235 Operations Report 4-MAY-2021

SOL: 2

Name of person filing report: Atila Meszaros

Non-nominal systems: Musk Observatory and Power System

Notes on non-nominal systems: Nothing to report

Generator: Working nominally. The battery is charging through the battery charger connected from the Science Dome.

Hours run: 0

From what time last night: the generator didn’t run

To what time this morning: the generator didn’t run

List any additional daytime hours when the generator was run: N/A

Solar— SOC 90% at 7:01 pm (Before generator is run at night)

Notes on power system: This morning the charge controllers were down, breakers jumped. We have been having energy through the day, but we went into power save mode during the morning. Shannon fixed the system in the morning and the SOC jumped again to 100% this afternoon. It has been steady since then.

Diesel Reading – 47%

Station Propane Reading – 80%

Outpost Propane Reading – 1. 72% 2. 80%

Water (loft tank): 25 gallons

Water Meter: 150994.0 units

Water (static tank): 400 gallons

Static to Loft Pump used – yes

Water in GreenHab: 0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: Yes

Sojourner rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Spirit rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Opportunity rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Curiosity rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Perseverance rover used: Yes

Hours: 208.

Beginning charge: 100%

Ending charge: 99%

Currently charging: Yes

Notes on rovers: None

ATV’s Used: Yes

Reason for use: EVA#2 and EVA#3

Oil Added? No

ATV Fuel Used: N/A

# Hours the ATVs were used today: 0

Notes on ATVs: ATV’s will be filled with Free Ethanol Gas to be ready for tomorrow’s EVA.

HabCar used and why, where?: Nothing to report

CrewCar used and why, where?: Nothing to report

General notes and comments: Nothing to report

Summary of internet: Working nominally.

Summary of suits and radios: All radios nominal, all functional suits currently charging.

Summary of Hab operations: We identified a small leak under the kitchen sink, but it was just a pipe connection that has already been connected and secure. It looks like there’s some water activity in the toilet room. The floor is getting constantly wet every time we use it. We caught a little rodent during the morning and it was released in Kissing Camels.

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Geology Class in the morning with Shannon.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Rotation system won’t work in the Musk Observatory. Waiting for instructions from Peter.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Thank you for being with us tonight!